OpenStudy (kayders1997):

g(p)=(p^2+3)(square root of p) what is the derivative?

OpenStudy (juan1857):

have you learned about the product rule?

OpenStudy (kayders1997):

Yes

OpenStudy (kayders1997):

I rewrote square root of p to p^1/2 also

OpenStudy (juan1857):



Nnesha (nnesha):

\[g(p) = (\sqrt{p^2}+3 )\] like this ?? or \[g(p)= \sqrt{p} +3 \] ??